# 3DGen-Bench: Comprehensive Benchmark Suite for 3D Generative Models

### v1

@@ 5.2 Human Preference Evaluations
Text-to-3D. According to our criteria, the top-3 methods in text-to-3D track are MVDream [61], LucidDreamer [32], and Magic3D [33], achieving average Elo scores of 1177.66, 1112.21, and 1088.93, respectively. The model ranking allows us to explore each model’s strengths and weaknesses and gain some insights for further improvement. The success of MVDream [61] demonstrates the advancement of multi-view diffusion model, which not only achieves the effective combination of 2D prior and 3D prior but also facilitates the information exchange between different views. Building on the 3DGS representation, LucidDreamer [32] performs well on geometry and texture quality. Magic3D [33] benefits from a ”two-stage” optimization strategy but is somehow short on texture quality.
Image-to-3D. On the image-to-3D track, Wonder3D [41], OpenLRM [17], and Stable Zero123 [39] emerge as the frontrunners, achieving average Elo scores of 1304.05, 1279.67, and 1200.69, respectively. Wonder3D [41] employs a multi-view cross-domain attention mechanism, which again verifies the significance of 2D&3D domain blending and multi-view consistency. Benefiting from Objarverse [8, 7] datasets, OpenLRM reinvigorates the transformer-based encoder-decoder framework. Meanwhile, Stable Zero123 [39] achieves a comprehensive improvement over Zero123-XL [39], but still leaves room for improvement in geometric details.

### v2

@@ 1 Introduction
In this paper, we propose the first comprehensive benchmark that incorporates large-scale human preferences and uniforms the evaluation of text-to-3D and image-to-3D generation, named 3DGen-Bench. Our main contributions can be summarized in three aspects: 1) Comprehensive model zoo for 3D generation. To build this dataset, we carefully curate over 1,000 text and image prompts across various categories, covering single-object, multi-object, and micro-scene, which is approximately tenfold the number of prompts designed in previous works [16, 77]. With the proposed prompt set, we evaluate 19 representative open-sourced generative models, including 9 for text-to-3D and 13 for image-to-3D, yielding over 10,000 3D assets and more than 58,000 battle pairs. 2) Large-scale multi-criteria human preference collection. To gather extensive human feedback, we build a voting platform open to community, dubbed as 3DGen-Arena, which adopt the anonymous pairwise battle manner, enabling users to express their preferences between pairs of 3D generative models. Considering the complexity of 3D contents, we identified five key dimensions for assessment [77]: Geometry Plausibility, Geometry Details, Texture Quality, Geometry-Texture Coherency, and Prompt-Asset Alignment. In practice, we collect feedback from two user groups: the general public and specialized experts. Although votes from public users are valuable for their volume, they may introduce noises that can affect subsequent tasks. To mitigate this, we also engage experts whose annotations offer a more reliable data source for human preferences. In addition to the pairwise comparisons from 3DGen-Arena, we also ask our experts to assign absolute scores for each 3D model, which is more suitable for large-scale evaluation. 3) Human-aligned automatic scoring models. The collected human preference dataset enables the training of a scoring model. Inspired by [50, 25], we finetune a pre-trained CLIP model [55] on our proposed dataset in a Reward Modeling(RM) style, named 3DGen-Score, which is designed to predict the human preference between two 3D models generated from the same prompt. Specifically, the model takes a prompt along with four-view normal and RGB renderings from two models as input and returns a win-rate tuple for five dimensions. Furthermore, leveraging the success of Multi-modal Large Language Models(MLLMs), we fine-tune LLaVA [35] to evaluate 3D models in a Question&Answer(QA) style, termed as 3DGen-Eval, which is encouraged to provide explanations to enhance credibility.
We have carried out comparative experiments on test data, evaluating our models against commonly used metrics such as CLIP-Score [18], Aesthetic score [59], GPTEval3D [77], and others. Experimental results demonstrate that our proposed models exhibit strong alignment with human references, comprehensively outperforming all other metrics in both pairwise alignment and ranking correlation. Meanwhile, each model offers distinct advantages. 3DGen-Score excels in human alignment, efficiency, and scalability, while 3DGen-Eval offers superior interpretability and robustness, matching or surpassing GPT-4V without additional costs. Furthermore, we also experimentally validate the potential of our scoring model in optimizing the quality of generated models served as a loss function or reward model. We believe that our work will play a significant role in advancing 3D evaluation, paving the way for high-quality 3D generation.
@@ 5.2 Human Preference Evaluations
Text-to-3D. According to our criteria, the top-3 methods in text-to-3D track are MVDream [61], LucidDreamer [32], and Magic3D [33], achieving average Elo scores of 1177.66, 1112.21, and 1088.93, respectively. These results shed light on the distinct strengths and weaknesses of each method, offering insights that can inform future improvements in 3D generation techniques. The success of MVDream [61] demonstrates the advancement of multi-view diffusion model, which not only achieves the effective combination of 2D prior and 3D prior but also facilitates the information exchange between different views. Building on the 3DGS representation, LucidDreamer [32] performs well on geometry and texture quality. Magic3D [33] benefits from a ”two-stage” optimization strategy, significantly enhancing the geometric plausibility of generated 3D models. However, its texture quality remains somewhat unsatisfactory, primarily due to the inherent limitations of SDS loss.
Image-to-3D. On the image-to-3D track, Wonder3D [41], OpenLRM [17], and Stable Zero123 [39] emerge as the frontrunners, achieving average Elo scores of 1304.05, 1279.67, and 1200.69, respectively. Wonder3D [41] employs a multi-view cross-domain attention mechanism, which again verifies the significance of 2D&3D domain blending and multi-view consistency. Benefiting from Objarverse [8, 7] datasets, OpenLRM reinvigorates the transformer-based encoder-decoder architecture, enhancing the quality and diversity of generated 3D objects. Meanwhile, Stable Zero123 [39] achieves a comprehensive improvement over Zero123-XL [39] in multiple dimensions, but still leaves room for improvement in geometric details.